# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v2

@@ Acknowledgements
       $run=0$  Randomly order nodes  for $v\in S$ do
             for $u\in V\setminus S$ do
                   $S^{\prime}=S$  Remove $v$ from $S^{\prime}$ and replace with $u$if $EI(S^{\prime})>EI(S)$ then
                         $S=S^{\prime}$$run=1$
@@
IM for large networks
To demonstrate the performance of $\mathsf{BOPIM}$ on large networks, we present the results for IM on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes. We fix $\lambda=0.05$ and $T=10$ while varying $k=25,50,\dots,100$. Because of the long run times, we set $N_{0}=5,\ B=20$ and only use 3 MC simulations for $\mathsf{BOPIM}$, but still use 25 for Random and Random Degree. The influence spread results and run times are in Figure 7. We can see that the proposed method provides comparable results to the greedy algorithm, but is faster for all but the first value of $k$. While $\mathsf{BOPIM}$ performs similarly to Random Degree, it greatly outperforms Dynamic Degree.
@@
Robustness experiments
@@
First, we are interested in the performance of the proposed method when the fitting steps are done on one value of $k^{\prime}$, but then we find the optimal seed set for $k\neq k^{\prime}$. Specifically, we let $k^{\prime}$ be fixed and then run the $\mathsf{BOPIM}$ algorithm for $B$ BO loop iterations as usual. After the last iteration, however, we re-fit the GP regression model and use the acquisition function to find the optimal seed set of size $k$ which is not necessarily the same as $k^{\prime}$. We then use this seed set to compute the number of influenced nodes on the network.
@@
In the following experiments, we fix $\lambda=0.05$ and $T=10$ and look at the Reality and Hospital networks. All other hyper-parameters are the same as in the previous experiments. We consider four different methods, Low, Mid, High and Mix. For Low, Mid and High, the BO routine is carried out setting $k^{\prime}=1,4$ and $7$, respectively. For Mix, we randomly sample $k^{\prime}$ from a uniform distribution on $\{1,2\dots,7\}$ at each stage in the initial fitting as well as the BO loop. For each method, we then compute the optimal seed set of size $k$ for $k=1,\dots,7$. The results are repeated 25 times and averaged for each value of $k$. We compare the results with the original $\mathsf{BOPIM}$ algorithm where the same $k$ is used in the training phase of the algorithm as well as the final influence spread evaluation.
@@
Note that even when the network is trained on seed set of size $k^{\prime}$ and then tested on sets of size $k=k^{\prime}$, this is still not equivalent to the original $\mathsf{BOPIM}$ algorithm. In the original algorithm, the model is fit on one value of $k$ throughout the $B$ BO iterations and then the set seed which corresponded to the largest influence spread is then chosen as the optimal set. In these robustness simulations, on the other hand, the final posterior distribution is used to select the optimal seed set. Therefore, Low, for example, is not equivalent to $\mathsf{BOPIM}$ even when $k^{\prime}=k=1$.
@@
Now, the results are in Figure 9. In the Reality network, for $k=2$, Low comes closest to the original $\mathsf{BOPIM}$ algorithm. When $k=3,4,5$, Mid performs well while High does the best for large $k$. Mix, however, yields seed sets with noticeably smaller influence spreads. In the Hospital network, Low, Mid and High all perform comparably well when $k\geq 4$ with Low and Mid yielding slightly larger optimal spreads for smaller values of $k$. Again, Mix is clearly performing quite poorly.
@@
These findings are promising for the performance of the $\mathsf{BOPIM}$ algorithm. In particular, it means that even if we run the algorithm for one value of $k^{\prime}$, the results can be used to select seed sets of various sizes. Additionally and not surprisingly, these results also show that as $|k^{\prime}-k|$ increases, the performance seems to decrease slightly. Thus, the method will yield better seed sets when $k^{\prime}\approx k$. This not only shows the generalizability of the method, but also can save computing time as the algorithm may not need to be re-run for each value of $k$. Finally, a single $k^{\prime}$ should be used in the algorithm’s model fitting as the Mix method consistently performed poorly.
@@
The results are in Figure 9. For the Reality network, both Mid and High perform similarly and nearly approximate the correctly-specified algorithm. Low, however, yields seed sets with smaller influence spreads. In the Hospital network, Low, Mid and High all perform comparably well in relation to the original algorithm.
Again, these results are favorable for the $\mathsf{BOPIM}$ algorithm. They show that the algorithm can be fit on almost any value of $T^{\prime}$ and the resulting seed set will yield large influence spreads for nearly all $T$. Moreover, they imply that the optimal seed set is not very sensitive to the number of snapshots used to aggregate the temporal network. For example, if a node is influential with $T=5$ snapshots, it is also likely to be influential with $T=10$ snapshots. This is somewhat expected as central nodes should be influential regardless of the network aggregation scheme. Thus, the proposed algorithm is relatively robust to the choice of $T$.

### v4

@@ 3. Greedy algorithm for $AEI(\cdot)$
       $run=0$  Randomly order nodes  for $i$ such that $x_{i}^{*}=1$ do
             for $j$ such that $x_{j}^{*}=0$ do
                   $\tilde{\bm{x}}=\bm{x}^{*};\ \tilde{x}_{i}=0;\ \tilde{x}_{j}=1$if $AEI(\tilde{\bm{x}})>AEI(\bm{x}^{*})$ then
                         $\bm{x}^{*}=\tilde{\bm{x}}$$run=1$
@@
4. Connection with Trust regions
The greedy algorithm can be seen as a relative of the Trust Region (TR) methodology, made popular in e.g., Eriksson et al., (2019); Wan et al., (2021); Papenmeier et al., (2023). Let $\bm{x}^{*}$ be the current best solution, i.e., seed set which has yielded the largest influence spread. Then $\mbox{TR}_{L}(\bm{x}^{*})$ is the Trust region of radius $L$ around $\bm{x}^{*}$, defined as all points within Hamming distance $L$ of $\bm{x}^{*}$, i.e.,
$\mbox{TR}_{L}(\bm{x}^{*})=\left\{\bm{x}\mid d_{H}(\bm{x},\bm{x}^{*})\leq L,\ \sum_{i=1}^{n}x_{i}=k\right\}.$
Of course for our problem, we can only consider points within the TR that also satisfy our cardinality constraint, which yields the second condition required for the TR.
Now, instead of searching the entire combinatorial input space for the optimal solution, a local search is performed within the TR. The radius $L$ is adjusted dynamically throughout the algorithm, increasing if a new optimum is found, and decreasing otherwise. When the radius falls below a certain threshold, the search may restart at a new initial condition.
With this in mind, we can see how the proposed greedy algorithm is similar to the TR framework. In the greedy algorithm, we have some current seed set $\bm{x}^{*}$ and we swap one seed node with an unseeded node to generate a candidate seed set $\tilde{\bm{x}}$. It is clear that
$d_{H}(\bm{x}^{*},\tilde{\bm{x}})=2,$
since the two seed sets exchange a single node. Note that two is also the shortest possible distance two eligible seed sets can be; a Hamming distance of one is impossible as it would imply different sized seed sets. Thus, the greedy algorithm can be considered as a modified TR search centered on $\bm{x}^{*}$ with $L=2$ fixed throughout the algorithm, albeit with a more restrictive search path.
5. IM for large networks
To demonstrate the performance of $\mathsf{BOPIM}$ on large networks, we present the results for IM on the DNC email network (Rossi and Ahmed,, 2015) with $n=1891$ nodes. We fix $\lambda=0.05$ and $T=10$ while varying $k=25,50,\dots,100$. Because of the long run times, we only consider $\mathsf{BOPIM-H}$ using 5 MC repetitions, while still using 25 for Random and Random Degree. The influence spread results and run times are in Figure 9. We can see that the proposed method provides comparable results to the greedy algorithm, but is significantly faster. While $\mathsf{BOPIM-H}$ performs similarly to Random Degree, it greatly outperforms Dynamic Degree.
@@ 5. IM for large networks
6. Surrogate function validation
Settings
In these simulations, we quantify the fit of the surrogate model to the observed values of the objective function. Towards this end, we first run Algorithm 1 with $N_{0}=5$ and $B=20$ and save the final posterior distribution of $\beta_{0}$. Using this posterior distribution, we then compute out-of-sample influence spread predictions on $N=100$ seed sets and compare with the observed values of the objective function. If the surrogate model fits well, then its predicted influence spread will be close to the true influence spread. We quantify this similarity using the mean absolute prediction error (MAPE). If $\hat{\beta}_{0}$ is the posterior median of $\beta_{0}$, and $\bm{x}^{*}$ corresponds to the new seed set, then
$\tilde{y}=\hat{\beta}_{0}+\kappa(\bm{x}^{*},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-\hat{\beta}_{0}{\bf 1}_{N_{0}+b})$
where $\kappa(\bm{x}^{*},{\bf X})_{j}=\kappa(\bm{x}^{*},\bm{x}_{j})$ for $j=1,\dots,N_{0}+B$ and ${\bf 1}_{n}$ is the vector of ones of length $n$. The MAPE is computed as
$\frac{1}{N}\sum_{i=1}^{N}|\tilde{y}_{i}-y_{i}|$
We stress that the goal of this experiment is to ensure that the surrogate model reasonably approximates the influence spreading process. Note that the observed values of the influence spread, $y_{i}$, are treated as the ground-truth, even though they were observed with noise from MC simulations. Thanks to the BO framework, we could obtain an entire predictive distribution, i.e. $p(\tilde{y}|{\bf X},\bm{y})$. While quantifying the uncertainty in our predictions of the influence spread is a by-product of this algorithm, it is not the primary goal of this section. Rather, we simply want to know if the predictive mean is similar to the observed influence spreads. Thus, MAPE is an appropriate metric of interest.
We compute the MAPE for each of the four data sets. We fix $\lambda=0.05,0.05,0.01,0.01$, and $k=3,4,6,10$ for the Reality, Hospital, Bluetooth and Conference 2 datasets, respectively. $k$ was chosen as approximately 5% of the total number of nodes for each network. We consider the proposed algorithm using both the Hamming ($\mathsf{BOPIM-H}$), as well as the Jaccard kernel ($\mathsf{BOPIM-J}$). As a baseline method, we also consider intercept-only predictions, where we simply take the mean influence spread of 25 seed sets randomly sampled proportional to a node’s degree.
For the out-of-sample predictions, we consider two node sampling schemes. One samples nodes at random while the other samples proportional to their degree on the aggregate network, as in $\mathsf{BOPIM}$. We expect the model to have a closer fit when the test points are sampled in the same way as the training points. Results are averaged over 25 Monte Carlo (MC) replications.
Results
The results are in Table 1. Note that the MAPE results are interpreted as follows: for the Hamming kernel on the Conference 2 network with degree-based sampling of seed sets, for example, the MAPE value of 1.06 means that, on average, the estimated influence spread of the surrogate function is within about one node of the true influence spread. We can see that for all networks and both sampling schemes, the MAPE values are quite small. We also notice very little difference in the results between the two kernels and intercept-only model. This is not too surprising as the $\mathsf{BOPIM}$ algorithm also uses an intercept-only mean function, and here we are making predictions about the mean. Of course, the intercept-only predictions cannot be used for seed selection as it gives no information on node importance.
Additionally, the degree-based sampling always leads to smaller MAPE values, which is to be expected since this is how the $\mathsf{BOPIM}$ algorithm samples points for the initial fitting stage. Indeed, the surrogate model fits the true objective function more closely when the seed nodes have a large degree, which is where we hypothesize that the global optimum lies. These results give use confidence that our surrogate function is doing an adequate job of modeling the true objective function.
Finally, while for all networks the Degree setting yields lower MAPE values, this difference is most pronounced for the Conference 2 network. This is likely due to the degree distributions of the various networks. In Figure 10, we report a histogram of the degree distribution for each network. For Conference 2, the vast majority of nodes have a very low degree. Since there are fewer nodes with large degree, when we sample proportionally to the node’s degree, it is more likely that the same nodes are being chosen. This increases the likelihood that the out-of-sample seed sets are similar to the seed sets that the model was already trained on, lowering the MAPE value. On the other hand, when the out-of-sample seed sets are randomly sampled, we will sample many nodes with very low degree, yielding seed sets unlike those that the model has been trained on, increasing the MAPE values.
Dataset Sampling Kernel MAPE (se)
Reality Random Intercept 3.77 (0.08)
Hamming 4.24 (0.23)
Jaccard 4.65 (0.22)
Degree Intercept 3.42 (0.06)
Hamming 3.09 (0.14)
Jaccard 3.37 (0.13)
Hospital Random Intercept 4.76 (0.09)
Hamming 4.58 (0.16)
Jaccard 4.61 (0.12)
Degree Intercept 3.05 (0.05)
Hamming 2.92 (0.07)
Jaccard 2.86 (0.08)
Bluetooth Random Intercept 2.49 (0.07)
Hamming 2.62 (0.09)
Jaccard 3.40 (0.16)
Degree Intercept 1.59 (0.02)
Hamming 1.56 (0.05)
Jaccard 2.02 (0.09)
Conference 2 Random Intercept 7.86 (0.07)
Hamming 6.87 (0.13)
Jaccard 5.06 (0.10)
Degree Intercept 1.01 (0.02)
Hamming 1.06 (0.04)
Jaccard 1.83 (0.11)
Figure 10: Degree distributions for the networks in this paper.
7. Initial sampling ablation study
In this section, we consider an alternative sampling scheme for the initial observations in the BO algorithm. We remark that nodes with high degrees on the temporally aggregated network, $\tilde{G}=\cup_{t}G_{t}$, are good candidates for the optimal seed set. Additionally, we would like to sample nodes that are “far apart” on the graph to ensure better influence spread. Therefore, instead of randomly sampling $\bm{x}_{i}$ from $\mathcal{X}$, we sample nodes proportional to a metric which accounts for degree and distance from the other nodes in the seed set. For the first node, we sample it proportional to the degree on the temporally aggregated network. Mathematically, this means sampling from a distribution with probability mass function $\mathsf{P}(Z=j)=d_{j}/\sum_{k}d_{k}$ where $d_{j}$ is the degree of node $j$ on $\tilde{G}$ for $Z\in\{1,\dots,n\}$. For the remaining $K-1$ nodes, we compute the average distance between the node $i$, and nodes already sampled for the seed set, call it $\nu_{i}^{(S)}$. We standardize both of these measures by converting to a $z$-score, i.e.,
$\tilde{d}_{i}=\frac{d_{i}-\bar{d}}{s_{d}},\ \ \tilde{\nu}_{i}=\frac{\nu_{i}-\bar{\nu}}{s_{\nu}}$
where $\bar{d}$ and $s_{d}$ are the sample mean and standard deviation of the degrees, $d_{1},\dots,d_{n}$, respectively, and similarly for $\bar{\nu}$ and $s_{\nu}$. Then we sample nodes proportionally to $\bm{p}_{\varrho}=(p_{1},\dots,p_{n})$ where
$p_{i}=\tilde{d}_{i}+\varrho\tilde{\nu}_{i}$
where $\varrho>0$ is a tuning parameter which controls between preferring nodes with high degree (small $\varrho$) with preferring nodes far apart (large $\varrho$). For all experiments, we set $\varrho=1$ for an equal balance. This sampling scheme ensures that the initial seed sets prioritize large degree nodes that are not nearby on the graph.
We compare this new sampling scheme (coined Space-degree) with the originally proposed sampling scheme (Degree-only) on the Reality network for $k=2,3,\dots,7$. For each value of $k$, we run the $\mathsf{BOPIM-H}$ algorithm, obtain the optimal seed set and compute the influence spread. This process is repeated 25 times and the average influence spread is reported in Table 2. The influence spreads of the two methods are almost identical, so we only present the simpler Degree-only sampling scheme in the main manuscript.
Sampler / $k$ 2 3 4 5 6 7
Degree-only 22.0 (0.2) 25.1 (0.1) 27.4 (0.1) 29.2 (0.1) 30.5 (0.1) 31.7 (0.1)
Space-degree 21.9 (0.3) 25.2 (0.1) 27.3 (0.2) 29.2 (0.1) 30.8 (0.1) 31.6 (0.1)
Table 2: Average influence spread (with standard error) for the degree-only and space-degree initial sampling schemes.
8. Robustness experiments
@@
First, we are interested in the performance of the proposed method when the fitting steps are done on one value of $k^{\prime}$, but then we find the optimal seed set for $k\neq k^{\prime}$. Specifically, we let $k^{\prime}$ be fixed and run $\mathsf{BOPIM-J}$ for $B$ BO loop iterations as usual. After the last iteration, however, we re-fit the GP regression model and use the acquisition function to find the optimal seed set of size $k$ which is not necessarily the same as $k^{\prime}$. We then use this seed set to compute the number of influenced nodes on the network.
@@
In the following experiments, we fix $\lambda=0.05$ and $T=10$ and look at the Reality and Hospital networks. All other hyper-parameters are the same as in the previous experiments. We consider four different methods, Low, Mid, High and Mix. For Low, Mid and High, the BO routine is carried out setting $k^{\prime}=2,4$ and $7$, respectively. For Mix, we randomly sample $k^{\prime}$ from a uniform distribution on $\{2,3,\dots,7\}$ at each stage in the initial fitting as well as the BO loop. For each method, we then compute the optimal seed set of size $k$ for $k=2,3,\dots,7$. The results are repeated 25 times and averaged for each value of $k$. We compare the results with the original $\mathsf{BOPIM}$ algorithm where the same $k$ is used in the training phase of the algorithm as well as the final influence spread evaluation.
@@
Note that even when the network is trained on seed set of size $k^{\prime}$ and then tested on sets of size $k=k^{\prime}$, this is still not equivalent to the original $\mathsf{BOPIM}$ algorithm. In the original algorithm, the model is fit on one value of $k$ throughout the $B$ BO iterations and then the set seed which corresponded to the largest influence spread is then chosen as the optimal set. In these robustness simulations, on the other hand, the final posterior distribution is used to select the optimal seed set. Therefore, Mid, for example, is not equivalent to $\mathsf{BOPIM}$ even when $k^{\prime}=k=4$.
@@
The results are in Figure 11. In the Reality network, High tends to consistently comes closest to the original $\mathsf{BOPIM}$ algorithm for all $k$. The relative performance of Low decreases as $k$ increases (for $k>2$), while Mid’s relative performance improves. Mix, however, yields seed sets with noticeably smaller influence spreads. The trends are similar in the Hospital network, where High and Mid consistently yields influence spreads closest to that of the original $\mathsf{BOPIM}$ algorithm, while Low’s performance is best for $k=3,4$. The performance of Mix is even worse relative to the other methods compared to the Reality network.
@@
These findings are promising for the performance of the $\mathsf{BOPIM}$ algorithm. In particular, it means that even if we run the algorithm for one value of $k^{\prime}$, the results can be used to select seed sets of various sizes. Additionally and not surprisingly, the general trends show that as $|k^{\prime}-k|$ increases, the performance seems to decrease slightly. Thus, the method will yield better seed sets when $k^{\prime}\approx k$. This not only shows the generalizability of the method, but also can save computing time as the algorithm may not need to be re-run for each value of $k$. Finally, a single $k^{\prime}$ should be used in the algorithm’s model fitting as the Mix method consistently performed poorly.
@@
The results are in Figure 12. For the both networks, Mid and High perform similarly and almost perfectly approximate the correctly-specified algorithm. Low, however, yields seed sets with much smaller influence spreads. Again, these results are favorable for the $\mathsf{BOPIM}$ algorithm. They show that the algorithm can be fit on almost any value of $T^{\prime}$ and the resulting seed set will yield large influence spreads for nearly all $T$. Moreover, they imply that the optimal seed set is not very sensitive to the number of snapshots used to aggregate the temporal network. For example, if a node is influential with $T=5$ snapshots, it is also likely to be influential with $T=10$ snapshots. This is somewhat expected as central nodes should be influential regardless of the network aggregation scheme. Thus, the proposed algorithm is relatively robust to the choice of $T$.
@@
9. Horseshoe prior Gibbs sampler
  1. 1.
    ${\bf S}=diag(\eta_{1}^{2}\tau^{2},\dots,\eta_{n}^{2}\tau^{2})$.
  2. 2.
    Sample $\sigma^{2}|\bm{\beta},\bm{\eta},\bm{\nu},\tau^{2},{\bf Y}\sim\mathsf{IG}(a_{1}+(N_{0}+n)/2,b_{1}+(({\bf Y}-{\bf X}\bm{\beta})^{T}{\bf K}^{-1}({\bf Y}-{\bf X}\bm{\beta})+\bm{\beta}^{T}{\bf S}^{-1}\bm{\beta})/2)$.
  3. 3.
    Sample $\eta_{j}^{2}|\bm{\beta},\bm{\nu},\tau^{2},\sigma^{2}\sim\mathsf{IG}(1,\nu_{j}^{-1}+\beta^{2}_{j}/(2\tau^{2}\sigma^{2})$, $j=1,\dots,n$
  4. 4.
    Sample $\tau^{2}|\bm{\beta},\bm{\eta},\sigma^{2},\xi\sim\mathsf{IG}((n+1)/2,\xi^{-1}+\sum_{k=1}^{n}\beta^{2}_{k}/(2\eta^{2}_{k}\sigma^{2}))$.
  5. 5.
    Sample $\nu_{j}|\eta_{j}\sim\mathsf{IG}(1,1+\eta_{j}^{-2})$ for $j=1,\dots,n$.
  6. 6.
    Sample $\xi|\tau^{2}\sim\mathsf{IG}(1,1+\tau^{-2})$.